Tickets at a park are $1.50 for kids and $4.00 for adults. On c certain day 2200 people entered the park, and the the admission fees collected totaled $5050. How many children and how many adults were admitted?
Answer by user_dude2008(1862)

x+y=2200 ------> y = 2200-x
1.5x+4y=5050


1.5x+4y=5050
15x+40y=50500
15x+40(2200-x)=50500
15x+88000-40x=50500
-25x+88000=50500
-25x=50500-88000
-25x=37500
x=1500 ------> 1500 kids

y=2200-x=2200-1500 = 700
y = 700 ----> 700 adults